The Cobb family decided to go roller skating for their reunion. There were a total of 44 people there and all went roller skating. If an adult admission was $5.00 and a child's admission was $3.00 and they paid a total of $168.00, how many children were at the reunion? thank you
Answer by mananth(12270)

The Cobb family decided to go roller skating for their reunion. There were a total of 44 people there and all went roller skating. If an adult admission was $5.00 and a child's admission was $3.00 and they paid a total of $168.00, how many children were at the reunion? thank you
...
let adults be x
children be y
...
x+y = 44
5x+3y=168
...
multiply first equation 1 by -5
-5x-5y=-220
add it to equation 2
-2y=-52
y=26 children
...
m.ananth@hotmail.ca